OpenStudy (wwe123):

solve this plz √(x^(2)-3x) = 4x^(2)- 12 x leave answer in square root

OpenStudy (anonymous):

Have you tried squaring both sides of the equation and calling x^2=s and solving the second order equation for s as you normaly do and taking the square root of s?

OpenStudy (anonymous):

oh forget it, there are terms of order 1 and 3 so it wouldn't work

OpenStudy (anonymous):

aha, found it. Call x^2-3x=y then the equation gets like this: \[\sqrt{y} = 4y\]Solve for y, then solve for x knowing y

OpenStudy (wwe123):

it means √(y)=4y-12

OpenStudy (anonymous):

No, only 4y because the right hand side is 4 times whats in the square root

OpenStudy (wwe123):

sorry question is √(x^(2)-3x) = 4x^(2)- 12x - 3

OpenStudy (anonymous):

Then it is \[\sqrt{y}=4y-3\]

OpenStudy (wwe123):

√(x^(2)-3x) = 4x^(2)- 12- 3x

OpenStudy (wwe123):

√(x^(2)-3x) = 4x^(2)- 12x - 3 (wrong)

OpenStudy (anonymous):

Well then I guess my sugestion does not help.

OpenStudy (anonymous):

Are you sure it is right? Because I tried putting it into mathematica and the exact solution was HUGE.